Use the unit circle to find the value of sin 3(pi)/2 and cos 3(pi)/2
1
Expert's answer
2019-07-17T08:59:23-0400

Solution.

(a)x2+y2=1

This is the equation of unit circle,where radius equals 1.

Let us consider the right triangle ABC.Hypotenuse AB is radius itself.As long as R=1,AB=1.

(b)

"\sin\alpha=\frac{opposite}{hypotenuse}"

"\sin\alpha=\frac{BC}{AB}"

BC is AB projection on y-axis,so that BC=y.

Then:

"\sin\alpha=\frac{y}{1}""\sin\alpha=y"

(c)

"\cos\alpha=\frac{adjacent}{hypotenuse}"

"\cos\alpha=\frac{AC}{AB}"

AC is AB projection on x-axis,so that AC=x.Then:


"\cos\alpha=x"

At the point where

"\alpha=\frac{3\pi}{2}""x=0"

According to this,

"\cos\frac{3\pi}{2}=0"

"y=-1"

And that is why

"\sin\frac{3\pi}{2}=-1"

Answer:

"\sin\frac{3\pi}{2}=-1"

"\cos\frac{3\pi}{2}=0"